OpenStudy (jkbo):

Determine whether the function is linear or quadratic. Identify the quadratic, linear, and constant terms. y= (x-5)(5x+4)-5x^2 A.linear function linear term: 15x constant term: –25 B.quadratic function quadratic term: 5x^2 linear term: 15x constant term: –25 C.linear function linear term: -21x constant term: –20 D.quadratic function quadratic term: -5x^2 linear term: -21x constant term: –20

OpenStudy (austinl):

Just multiply it out! \[y= (x-5)(5x+4)-5x^2\] \[(x-5)(5x+4) = ?\]

OpenStudy (jkbo):

5x^2 -21x - 20 @austinL

OpenStudy (austinl):

Correct. So now we have \[(5x^2 -21x - 20)-5x^2\]

OpenStudy (jkbo):

what would I do from here?

OpenStudy (jkbo):

@austinL

OpenStudy (austinl):

You just simplify. You have only two terms with a factor of 2 so \[5x^2−21x−20−5x^2=5x^2−5x^2−21x−20\]

OpenStudy (jkbo):

I got -21x-20 = -21x-20

OpenStudy (jkbo):

@austinL

OpenStudy (austinl):

Now is that one of your answers? *hint, it is* *cough, cough C, cough*

OpenStudy (jkbo):

hahah Thanks man.

OpenStudy (austinl):

Yep, no problem!

OpenStudy (jkbo):

@austinL can you help me with one more its easy

OpenStudy (austinl):

Sure, I can try.

OpenStudy (jkbo):

Identify the vertex and the y-intercept of the graph of the function y= -3(x+2)^2+5 A.vertex: (2, -5); y-intercept: -12 B.vertex: (-2, -5); y-intercept: 9 C.vertex: (-2, 5); y-intercept: -7 D.vertex: (2, 5); y-intercept: -7

OpenStudy (jkbo):

@austinL

OpenStudy (austinl):

(x+2)(x+2)=?

OpenStudy (jkbo):

x^2 + 4x+4

OpenStudy (austinl):

Correct, now multiply that all by 3.

OpenStudy (jkbo):

3x + 12x + 12

OpenStudy (austinl):

now add 5 to that.

OpenStudy (jkbo):

to all of it?

OpenStudy (austinl):

No, just to the terms without a variable. :)

OpenStudy (jkbo):

3x+ 12x + 17

OpenStudy (austinl):

Wait, rewind. Multiply it by (-3) sorry.

OpenStudy (jkbo):

-3x-12x-17

OpenStudy (austinl):

then add 5 to that :D

OpenStudy (jkbo):

-3x-12x-12

OpenStudy (austinl):

\[3x^2-12x-12\]. So now what is your y intercept?

OpenStudy (jkbo):

-12

OpenStudy (austinl):

Right, is that in one of your answers?

OpenStudy (jkbo):

yes for A.

OpenStudy (austinl):

there you go!

OpenStudy (jkbo):

thank you

OpenStudy (austinl):

No problem!